OpenStudy (anonymous):

5=y/3 - 9 (Show Work) ( / = Over) HELP!

OpenStudy (anonymous):

y/3 = 14 y = 42.

OpenStudy (saifoo.khan):

\[5 + 9 = \frac y3\] \[14 = \frac y3\] \[y = 3 \times 14\] \[y = Answer\]

OpenStudy (anonymous):

muahah. got it first saifoo ;)

OpenStudy (saifoo.khan):

I am happy that you won! :)